OpenStudy (anonymous):

The vertex of f(x) = 2(x-2)2 +10 is (-2, 10). True False

OpenStudy (michele_laino):

I think true

OpenStudy (michele_laino):

It's true because, the x coordinate of the vertex is: \[\frac{ -b }{ 2a }=-\frac{ -8 }{ 2*2 }=2\] now, if you substitute x=2 into the equation of your parabola, you would get the y-coordinate of the vertex, which is:y=f(2)=10

OpenStudy (michele_laino):

Sorry the vertex is (2,10)

OpenStudy (michele_laino):

So your answer is false, sorry again!